An airplane took 6 hours to fly between two cities, flying directly into a 50 mph wind. The return trip with the same wind took only 4 hours. What is the distance between the cities?
Answer by mananth(16949)

let speed in still air = x
wind speed 50
..
against wind speed x- 50
with wind speed 50
..
against wind 6 hours
with wind 4 hours
..
6(x-50)=4(x+50)
6x-300 =4x+200
6x-4x=300+200
2x=500
/ 2
x = 250
Distance = (250 +50)* 4
Distance = 1200 miles